# Pruning the Paradox: How CLIP’s Most Informative Heads Enhance Performance While Amplifying Bias

### v2

@@ 4.3 High CCS heads are crucial for out-of-domain (OOD) detection
Previously, we established that high-CCS heads in CLIP models are crucial for image and video tasks and pruning them leads to significant drop in performance. Now, we investigate if these high CCS heads learn spurious features leading to social biases. For this, we perform soft pruning experiment on FairFace and SocialCounterFactuals datasets. Here given neutral text prompts of 104 occupations as show in tables  5 and  6, we measure MaxSkew across race and gender in the datasets. Tables  7 and  8 show the results on FairFace and SocialCounterFactuals datasets respectively.
On the FairFace dataset, pruning high-CCS heads consistently reduces the MaxSkew values for both race and gender across all models. For example, in the ViT-B-16-OpenAI model, pruning high-CCS heads drops the race MaxSkew from 64.61 to 55.55 and the gender MaxSkew from 40.19 to 38.21. Similar reductions are observed across all ViT-B and ViT-L variants. These drops, although modest in some cases, indicate a consistent trend: high-CCS heads are contributing disproportionately to skewed model predictions. The effect is even more evident on the SocialCounterfactuals dataset, where MaxSkew values drop substantially upon pruning high-CCS heads. For instance, in ViT-B-32-OpenAI, the gender MaxSkew falls from 4.05 to 1.22, and race MaxSkew from 3.65 to 2.43. Similar reductions occur for other ViT variants, with some pruned models showing more than 50% decrease in bias.
These results reveal a fundamental paradox at the heart of CLIP models: high-CCS heads, while critical for strong performance in tasks such as classification, retrieval, and concept alignment, are also the primary contributors to social bias. This paradox emerges from CLIP’s contrastive learning objective, which optimizes alignment between images and their paired text across large, uncurated datasets. In doing so, the model often absorbs and amplifies spurious correlations between visual features and demographic attributes. High-CCS heads, by virtue of their consistent focus on semantically aligned regions, become particularly susceptible to reinforcing these correlations. Pruning these heads leads to a notable reduction in model bias, as shown in our experiments, but also comes at the cost of reduced performance—a clear trade-off between fairness and utility. This performance-bias paradox underscores the complex role of high-CCS heads: they are both enablers of semantic understanding and carriers of learned stereotypes. The CCS metric, in this context, provides a valuable lens for navigating this tension. It not only aids in interpreting model behavior but also offers a lightweight intervention—soft-pruning—that mitigates bias without requiring expensive fine-tuning.

### v3

@@ 4.3 High CCS heads are crucial for out-of-domain (OOD) detection
Previously, we established that high-CCS heads in CLIP models are crucial for image and video tasks and pruning them leads to significant drop in performance. Now, we investigate if these high CCS heads learn spurious features leading to social biases. For this, we perform soft pruning experiment on FairFace and SocialCounterFactuals datasets. Here given neutral text prompts of 104 occupations***List of occupations and prompts can be foudn in Appendix, we measure MaxSkew across race and gender in the datasets. Tables  8 and  9 show the results on FairFace and SocialCounterFactuals datasets respectively.
On the FairFace dataset, pruning high-CCS heads consistently reduces the MaxSkew values for both race and gender across all models. For example, in the ViT-B-16-OpenAI model, pruning high-CCS heads drops the race MaxSkew from 64.6 to 55.6 and the gender MaxSkew from 40.2 to 38.2. Similar reductions are observed across all ViT-B and ViT-L variants. These drops, although modest in some cases, indicate a consistent trend: high-CCS heads are contributing disproportionately to skewed model predictions. The effect is even more evident on the SocialCounterfactuals dataset, where MaxSkew values drop substantially upon pruning high-CCS heads. For instance, in ViT-B-32-OpenAI, the race MaxSkew falls from 4.1 to 1.2, and gender MaxSkew from 3.7 to 2.4. Similar reductions occur for other ViT variants, with some pruned models showing more than 50% decrease in bias.
These results reveal a fundamental paradox at the heart of CLIP models: high-CCS heads, while critical for strong performance in tasks such as classification, retrieval, and concept alignment, are also the primary contributors to social bias. Pruning these heads leads to a notable reduction in model bias, as shown in our experiments, but also comes at the cost of reduced performance—a clear trade-off between fairness and utility. This performance-bias paradox underscores the complex role of high-CCS heads: they are both enablers of semantic understanding and carriers of learned stereotypes. The CCS metric, in this context, provides a valuable lens for navigating this tension. It not only aids in interpreting model behavior but also offers a lightweight intervention—soft-pruning—that mitigates bias without requiring expensive fine-tuning.
@@ 5 Conclusion
6 Limitations
In this work, we experimented primarily on CLIP models. Although CCS metric established the fundamental paradox of performance vs social biases we haven’t proved for other vision language models. Hence, we leave extending for more vision language models for future work. Another limitation is the use of LLM models for concept labelling and judging which requires robust manual verification to limit any inconsistencies. Hence, scaling our work to much bigger models with more layers and heads can be a limitation.
@@ 6 Limitations
Model Concentrated Concept Ratio (CCR)
ViT-B-32-OpenAI 0.250
ViT-B-32-datacomp 0.273
ViT-B-16-OpenAI 0.200
ViT-B-16-LAION 0.231
ViT-L-14-OpenAI 0.125
ViT-B-14-LAION 0.143
Table 11: Concentrated Concept Ratio (CCR) across different CLIP model variants. Lower CCR in larger models (e.g., ViT-L-14) suggests more distributed concept representations.
Appendix B Redundancy and Functional Duplication Analysis.
Previously in section  4.1, we observed that larger models degrade less when high-CCS (Concept Contribution Score) attention heads are pruned. This raises the question of whether larger models exhibit greater redundancy—i.e., whether multiple heads are attending to the same concepts, providing robustness against pruning.
To quantify this, we introduce the Concentrated Concept Ratio (CCR), a metric that captures the extent to which concepts are redundantly represented across multiple high-CCS heads.
$CCR=\frac{C_{multi}}{H_{high}}$ (1)
where $H_{high}$ be the set of attention heads with high CCS, $C_{multi}$ be the number of unique concepts that are concentrated on more than one head within $H_{high}$. A higher CCR implies greater redundancy—i.e., the same concepts are captured by multiple heads. Conversely, a lower CCR suggests a more distributed and unique representation of concepts across heads.
Table  11 presents CCR values across models of varying sizes. Interestingly, larger models (e.g., ViT-L variants) show lower CCR, indicating that they tend to distribute conceptual information more evenly across heads. This aligns with the empirical finding that these models are more resilient to pruning of high-CCS heads.
To further examine whether redundancy exists across layers (functional duplication), we perform a layer-wise CCR analysis restricted to the last four transformer layers of each model. This focus is motivated by prior work  Gandelsman et al. , which shows that only the final four layers significantly affect model outputs. CCR is computed independently for each of the last four layers. However, the analysis reveals no consistent trend indicating that deeper layers exhibit greater concept duplication. This suggests that redundancy across layers is not a dominant factor and that the observed robustness in larger models is more likely due to distributed head-level representations rather than functional duplication across layers.
@@
Occupations
Prompt Example
A <occupation> A biologist
A photo of <occupation> A photo of biologist
A picture of <occupation> A picture of biologist
An image of <occupation> An image of biologist